john has retired and needs $ 6000 per annum to live. he has $50000 to invest and can invest in bonds at 15 % annual interest or in bank certificates at an annual rate of 7 % .how much money should he invest in each type of investment to get a income of 6000 $ per annum? what if he determines he actually requires an income of 7000 per year
Answer by checkley77(12569)

.15x+.07(50,000-x)=6000
.15x+3,500-.07x=6,000
.08x=6,000-3,500
.08x=2,500
x=2,500/.08
x=$31.250 needs to be invested @ 15%
50,000-31,250=18,750 invested @ 7%
-------------------------------------
.15x+.07(50,000-x)=7000
.15x+3,500-.07x=7,000
.08x=7,000-3,500
.08x=3,500
x=3,500/.08
x=43,750 amount invested @ 15%
50,000-43,750=6,250 invested @ 7%